# Interpretable representation learning of quantum data enabled by probabilistic variational autoencoders

### v2

@@ I Introduction
The rapid progress of classical machine learning (ML) is profoundly impacting our ability to analyze quantum systems and process quantum data, paving the way for new paradigms in scientific discovery [1]. In the quantum domain, by leveraging both supervised and unsupervised learning methods, researchers are now able to identify quantum phase transitions [2, 3, 4], model complex entanglement features [5], and improve quantum state tomography [6]. Notably, neural networks have been shown to offer compact and expressive representations of quantum states, capable of capturing their intricate correlations and phase structure [7].
@@
Beyond its predictive capacity, ML techniques with enhanced interpretability can serve as a lens through which the intrinsic properties of quantum systems can be unravelled [8]. For example, support vector machines have been employed to identify order parameters [9], symbolic regression methods have revealed underlying equations of motion [10], and neural architectures leveraging correlation functions have been used to map out phase diagrams of experimental quantum many-body systems [11]. Among these interpretable models, variational autoencoders (VAEs) [12] have attracted attention due to their ability to autonomously extract meaningful and minimal representations of some input dataset in their latent space [13]. When applied to physics datasets, this relates to extracting the meaningful physical features of the system of study [1]. In the quantum realm, VAEs have been used for instance to study phase transitions in spin systems [14, 15].
@@
Nonetheless, one of the key challenges when applying machine learning methods to quantum data lies in its intrinsically stochastic nature. As a result, physical observables must be inferred statistically over repeated measurements. This is particularly challenging in experimental platforms such as ultra-cold atoms in optical lattices [16] or trapped-ion quantum simulators [17], where performing repetitive measurement is costly, often leading to poor statistics. This fundamental stochasticity has motivated the development of ML approaches that operate directly on individual measurement outcomes, bypassing the need for full state reconstruction and instead learning meaningful features from single or ensembles of quantum snapshots [18, 4, 3, 19].
@@
In this sense, accounting for the intrinsic stochasticity of quantum data is essential for developing interpretable machine learning models. Since the physical properties of a quantum system are fundamentally encoded in its probability distribution, capturing this distribution is necessary for any meaningful characterization. This is particularly critical for VAEs: if they cannot reproduce the properties of the underlying probabilistic distribution of its input, it will not be able to create meaningful representations in its latent space. However, existing VAE applications on quantum data have not considered this stochasticity [14, 20, 21, 15, 22]. In particular, they commonly use of deterministic losses, which fundamentally misaligns with the goal of learning the probability distribution of the target quantum states. Such losses direct the model to capture only one-body, or mean-field, features, thereby neglecting nontrivial many-body correlations. It is hence critical to develop architectures that can cope with the stochastic properties of the quantum data, if one aims to use this method to extract physical features.
@@
To tackle these challenges, we propose two key adaptions to the VAE architecture. First, we train the model to directly approximate the underlying probability distribution of the input data, rather than exactly reconstructing particular samples. Second, we adapt the VAE’s decoder to be able to capture the stochastic properties of effectively any target quantum state. To achieve the latter, we draw inspiration from Neural Quantum States (NQS) [7], neural network architectures used as expressive variational ansätze capable of representing quantum probability distributions.
@@
To demonstrate the impact of the proposed improvements into the VAE framework for quantum data, we compare its resulting performance to that of a standard deterministic VAE on two paradigmatic spin models: the transverse-field Ising model with next-nearest neighbour interactions (NNN-TFIM) and long-range interactions (LR-TFIM). Through these models, we will explore and identify regimes where conventional VAEs struggle while the presented method provides physically interpretable representations. We further validate the approach on experimental snapshots of Rydberg atom arrays [16]. Without access to any prior labeling, knowledge of the governing Hamiltonian nor the number of relevant order parameters, the model is able to autonomously uncover the underlying phase structure, underscoring its potential as a powerful unsupervised and interpretable tool for the study of quantum many-body systems.
@@ II Methods
In this section we provide a detailed introduction to variational autoencoders (VAEs), outlining their fundamental principles. We then introduce the key adaptations required to effectively handle the stochastic nature of quantum data.
@@ II.1 Background on Variational Autoencoders
Variational Autoencoders (VAEs) consist of three main components: an encoder, a latent space, and a decoder [12]. The encoder compresses the input data into a lower-dimensional representation, namely the latent space. In contrast to autoencoders, the latent space of VAEs is parameterized by a probability distribution, typically defined as a multivariate Gaussian distribution. In practice, the encoder outputs the mean $\mu_{i}$ and variance $\sigma_{i}$ of the latent variables from which the values for each latent neuron $z_{i}$ are sampled and subsequently passed to the decoder (see Fig. 1). The decoder then reconstructs the input data based on the sampled latent variables.
@@
The loss function introduced in Eq. 1 contains two terms. First, a reconstruction loss, which enforces similarity between the input and output data. Second, a regularization term, which minimizes the Kullback-Leibler (KL) divergence between the latent space distribution $q(z_{i})\sim\mathcal{N}(\mu_{i},\sigma_{i})$ and the prior $p(z)$, typically defined as standard normal distribution $p(z)\sim\mathcal{N}(0,1)$. This term acts as a regularizer, enforcing the latent representation to align with the prior distribution. However, if all latent variables were to collapse to the prior, the decoder would lack sufficient information to faithfully reconstruct the data, leading to an increase in the reconstruction loss. This trade-off induces a natural competition between the two terms in the loss function, such that only the minimal set of latent neurons needed to properly reconstruct the input diverge from the prior (commonly referred to as an active neuron), while the other neurons are ”noised out” and collapse to $p(z)$ (referred to as passive neurons). Moreover, one can further scale the regularization term with a hyperparameter $\beta$ [13], such that the two terms are properly balanced. In practice, this mechanism forces the model to retain only the essential information from the input data in the latent space. When such data is associated with a few hidden variables, e.g. the degrees of freedom of a physical system, the resulting latent representation tends to align with these [1, 23, 24].
II.2 VAE with a generative decoder
@@
Figure 1: Schematic representation of the cpVAE. A spin configuration is input into the encoder neural network ENC. The latter outputs the means $\mu_{i}$ and the variances $\sigma_{i}$ parameterizing the Gaussian distribution from which the latent variables $z_{i}$ are sampled. Then, the decoder neural network DEC takes as input the latent variables and outputs the conditional probabilities of each spin be in a given state (up or down in the present scheme). During training, the decoder also receives as input the spin configuration to be reconstructed by the conditional probabilities, as indicated with the dotted arrow. Once trained, the latent neurons $z_{i}$ encode the main physical features of the physical system, which directly relate to its phase space. Moreover, one can fix the value of the latent variables and generate new configurations by autoregressively sampling the output conditional probabilities.
@@
As discussed in the introduction, prior applications of VAEs to quantum data have largely employed architectures with deterministic decoders [14, 20, 21, 15, 22]. In these models, the decoder is trained to reconstruct individual spin configurations using a deterministic reconstruction loss between input and output spin configurations, as for instance the mean squared error (MSE). This approach implicitly assumes a deterministic relationship between the latent representation and the output configuration, neglecting the intrinsic probabilistic nature of quantum measurement outcomes. Moreover, the use of a deterministic reconstruction metric enforces the model to learn average configurations across samples rather than capturing the full probability distribution of outcomes. This leads to a latent space that reflects only the mean-field structure of the data, failing to capture any probabilistic feature, as for instance correlations between nearby spins. In what follows, we refer to this baseline model as the deterministic VAE (dVAE) to distinguish it from the stochastic decoder architecture introduced in this work.
@@
Motivated by the recent work on VAEs for stochastic processes [24] and arguments on using generative classifiers for phase-classification [25], we present here the key changes that will allow VAE to correctly create interpretable representations of quantum data. First, we construct a decoder inspired by the recent advances in Neural Quantum States (NQS) [26]. For example, NQS have recently demonstrated impressive performance in ground-state searches and non-equilibrium dynamics [27]. By integrating an NQS-based decoder into the VAE framework, we ensure that the model can accurately capture and regenerate the statistical structure of the input quantum states.
@@
To account for the two previous arguments, we propose the use of an auto-regressive neural network (ARNN) for the decoder (see Fig. 1 and Appendix A for further details), and train it to model the conditional probabilities of each spin configuration, given the preceding spins. More precisely, the configuration of spin $i$ is given by the Bernouilli distribution. Its probability mass function $f$ over the two possible outcomes $x_{i}\in{1,-1}$ is
@@
To compute the KL regularization (second term in Eq. 1), we adopt the decomposition introduced in the Total Correlation Variational Autoencoder (TC-VAE) framework [28]. This approach enables a fine control over the structure of the latent space, by explicitly penalizing statistical dependencies between latent variables. This enforces the distinct degrees of freedom of some input physical data to be stored in independent latent neurons. The ELBO TC-decomposition consists in rewriting the KL regularization as
@@
In the latter, the first term on the right-hand side is the mutual information $I_{q}(z;x)$. The second term, referred to as the total correlation (TC), encourages the model to learn as so-called disentangled representations [28], where individual latent units capture distinct, minimally correlated physical features of the data. The third term is the dimension-wise KL which aligns the latent variables with the prior $\mathcal{N}(0,1)$. The hyperparameters, $\alpha$, $\beta$ and $\gamma$ are introduced to control the weight of each of these terms. The full loss with which the model is trained is thus
@@
$-\alpha I_{q}(z;n)-\beta\mathrm{KL}(q(z)||\prod_{j}q(z_{j}))$ (7)
@@ III Results
The first spin model is the transverse-field Ising model (TFIM) with next-nearest-neighbor (NNN) interaction. For weak NNN coupling, the model effectively reduces to the conventional TFIM, a setting where prior studies have applied deterministic VAEs [14, 20, 21]. In this regime, we show that this success is largely due to the fact that the relevant order parameter is a mean-field observable. As the strength of the NNN interaction increases, the system deviates from such mean-field behavior, and we observe that the deterministic VAE fails to reconstruct physical features, in contrast to the probabilistic one, effectively hindering the representation of the system in their latent space.
@@
The code needed to reproduce the results presented in each figure is available in the accompanying code repository [29].
@@ III.1.1 NNN-TFIM
We start by considering the Transverse Field Ising Model with nearest neighbour and next-nearest-neighbour interaction (NNN-TFIM) in a one dimensional chain, whose Hamiltonian has the form[30]
@@
with $J_{2}\in[0,1]$ and $X^{i},Z^{i}$ being the Pauli’s operators on spin $i$. The system under consideration has a size of $N=20$ with periodic boundary conditions.
The phase diagram of the NNN-TFIM can be characterized using two observables: the next-nearest-neighbor correlator $\langle Z^{i}Z^{i+2}\rangle$ and the magnetization $\langle Z^{i}\rangle$, as shown in Fig. 2a. For coupling strengths $J_{2}<0.5$, the nearest-neighbor interaction dominates, and the system effectively reduces to the conventional TFIM, a regime where deterministic VAEs had previous success [14, 20, 21].
@@
We then train the cpVAE with configurations arising from different points in the phase space of this spin model. During training, we see that a single neuron remains active, while all other neurons collapse to the prior and are noise out (see Appendix B for further details). This is to be expected since, as shown in Fig. 2a, the whole phase space can be characterized only by the second nearest-neighbour correlator. To mitigate the effect of hyperparameter tuning and to make the comparison as fair as possible, we then train the dVAE without KL regularization and with a latent dimension of 1. We note that not doing so, and considering a larger latent space, would allow the dVAE to store some stochastic features of the input sample in the latent space, for example, particular spins being up or down.
@@
We then benchmark the generative capabilities of both models. To do so, we first input configurations from the whole phase space to the encoder, extract the corresponding latent space, and then sample new configurations from the decoder based on the extracted latent variables. In Fig. 2(b), we show the $\langle Z^{i}Z^{i+2}\rangle$ correlator for the output configurations of the dVAE across the phase. While the model accurately reproduces the expected correlations in both the ferromagnetic and paramagnetic regimes, it fails to capture the antiferromagnetic phase. Importantly, this regime cannot be characterized by the magnetization only, the usual order parameter extracted in previous works. Since the alternating bond structure inherent to the frustrated ground state falls beyond any mean-field parameter, the deterministic decoder is unable to correctly reproduce this statistics, and hence the antiferromagnetic phase will never appear in the latent representation of a dVAE.

### v3

@@ I Introduction
The rapid progress of classical machine learning (ML) is profoundly impacting our ability to analyze quantum systems and process quantum data, paving the way for new paradigms in scientific discovery [iten2020dicvovering]. In the quantum domain, by leveraging both supervised and unsupervised learning methods, researchers are now able to identify quantum phase transitions [carrasquilla2017machine, cybinski2024speak, miles2023machine, huembeli2018identifying], model complex entanglement features [linda2024predicting], and improve quantum state tomography [torlai2018neural]. Generative ML in particular, has become an impactful tool in quantum physics, from the use of generative adversarial networks (GANs) for Hamiltonian learning  [kim2024hamiltonian] to diffusion models for quantum circuit synthesis [furrutter2024quantum]. However, such models are typically black boxes, hindering our ability to interpret the origin or reasons of their predictions.
@@
In this sense, ML techniques with enhanced interpretability can serve as a lens through which the intrinsic properties of quantum systems can be unraveled [wetzel2025interpretable]. For example, support vector machines have been employed to identify order parameters [ponte2017kernel], symbolic regression methods have revealed underlying equations of motion [kharkov2021discovering], and neural architectures leveraging correlation functions have been used to map out phase diagrams of experimental quantum many-body systems [miles2021correlator]. Among these interpretable models, variational autoencoders (VAEs) [kingma2013vae] have attracted attention due to their ability to autonomously extract meaningful and minimal representations of some input dataset in their latent space [higgins2017betavae]. When applied to physics datasets, this relates to extracting the meaningful physical features of the system of study [iten2020dicvovering]. In the quantum realm, VAEs have been used for instance to study phase transitions in spin systems [walker2020deep, naravane2023semi].
@@
Nonetheless, one of the key challenges when applying machine learning methods to quantum data lies in its intrinsically stochastic nature. As a result, physical observables must be inferred statistically over repeated measurements. This is particularly challenging in experimental platforms such as ultra-cold atoms in optical lattices [ebadi2021quantum] or trapped-ion quantum simulators [li2023probing], where performing repetitive measurement is costly, often leading to poor statistics. This fundamental stochasticity has motivated the development of ML approaches that operate directly on individual measurement outcomes, bypassing the need for full state reconstruction and instead learning meaningful features from single or ensembles of quantum snapshots [patel2022unsupervised, miles2023machine, cybinski2024speak, suresh2025interpretable].
@@
In this sense, accounting for the intrinsic stochasticity of quantum data is essential for developing interpretable machine learning models. Since the physical properties of a quantum system are fundamentally encoded in its probability distribution, capturing this distribution is necessary for any meaningful characterization. This is particularly critical for VAEs: if they cannot reproduce the properties of the underlying probabilistic distribution of its input, it will not be able to create meaningful representations in its latent space. However, existing VAE applications on quantum data have not considered this stochasticity [walker2020deep, wetzel2017unsupervised, jang2024unsupervised, naravane2023semi, frohnert2024explainable]. In particular, they commonly use deterministic losses, which fundamentally misaligns with the goal of learning the probability distribution of the target quantum states. Such losses direct the model to capture only one-body, or mean-field, features, thereby neglecting nontrivial many-body correlations. It is hence critical to develop architectures that can cope with the stochastic properties of the quantum data, if one aims to use this method to extract physical features.
@@
To tackle these challenges, we propose two key adaptions to the VAE architecture. First, we train the model to directly approximate the underlying probability distribution of the input data, rather than exactly reconstructing particular samples. Second, we adapt the VAE’s decoder to be able to capture the stochastic properties of effectively any target quantum state. To achieve the latter, we draw inspiration from Neural Quantum States (NQS) [carleo2017solving], neural network architectures used as expressive variational ansätze capable of representing quantum probability distributions (see Ref. [lange2024from] for an in depth review on the topic).
@@
To demonstrate the impact of the proposed improvements into the VAE framework for quantum data, we compare its resulting performance to that of a standard deterministic VAE on two paradigmatic spin models: the transverse-field Ising model with next-nearest neighbor interactions (NNN-TFIM) and long-range interactions (LR-TFIM). Through these models, we will explore and identify regimes where conventional VAEs struggle while the presented method provides physically interpretable representations. We further validate the approach on experimental snapshots of Rydberg atom arrays [ebadi2021quantum]. Without access to any prior labeling, knowledge of the governing Hamiltonian nor the number of relevant order parameters, the model is able to autonomously uncover the underlying phase structure, underscoring its potential as a powerful unsupervised and interpretable tool for the study of quantum many-body systems.
@@ II Methods
In this section, we provide a detailed description of the proposed pipeline, schematically presented in Fig. 1. We consider here a quantum system that can be tuned through two controllable parameters $\theta_{1}$, $\theta_{2}$ (see Fig. 1a). A measurement device records snapshots for a wide range of parameter values. This dataset of snapshots, which contains no information about the parameters $\theta_{1,2}$, is then used to train a VAE. As we will explain below, the VAE learns a compact, physically meaningful representation of the system’s physical descriptors. This representation can be accessed by inspecting the value of the latent neurons z for snapshots across the parameter space (see Fig. 1b). As we will show in  Section III, such latent representations will directly correlate with known order parameters of the systems under study, allowing us to identify different phases or regions throughout the parameter space. In the following, we outline the fundamental principles of VAEs, as well as the key adaptations required to effectively handle the stochastic nature of quantum data.
@@ II.1 Background on Variational Autoencoders
Variational Autoencoders (VAEs) consist of three main components: an encoder, a latent space, and a decoder [kingma2013vae]. The encoder compresses the input data into a lower-dimensional representation, namely the latent space. In contrast to autoencoders, the latent space of VAEs is parameterized by a probability distribution, typically defined as a multivariate Gaussian distribution. In practice, the encoder outputs the mean $\mu_{i}$ and variance $\sigma_{i}$ of the latent variables from which the values for each latent neuron $z_{i}$ are sampled and subsequently passed to the decoder (see Fig. 1c). The decoder then reconstructs the input data based on the sampled latent variables.
@@
The loss function introduced in Eq. 1 contains two terms. First, a reconstruction loss, which enforces similarity between the input and output data. Second, a regularization term, which minimizes the Kullback-Leibler (KL) divergence between the latent space distribution $q(z_{i})\sim\mathcal{N}(\mu_{i},\sigma_{i})$ and the prior $p(z)$, typically defined as standard normal distribution $p(z)\sim\mathcal{N}(0,1)$. This term acts as a regularizer, enforcing the latent representation to align with the prior distribution. However, if all latent variables were to collapse to the prior, the decoder would lack sufficient information to faithfully reconstruct the data, leading to an increase in the reconstruction loss.
This trade-off induces a natural competition between the two terms in the loss function, such that only the minimal set of latent neurons needed to properly reconstruct the input diverge from the prior (commonly referred to as an active neuron), while the other neurons are ”noised out” and collapse to $p(z)$ (referred to as passive neurons). Moreover, one can further scale the regularization term with a hyperparameter $\beta$ [higgins2017betavae], such that the two terms are properly balanced.
In practice, this mechanism forces the model to retain only the essential information from the input data in the latent space. This means that the VAE will noise out unnecessary latent neurons until finding a minimal set able to describe such input data. When the latter is associated with a few hidden variables, e.g. the degrees of freedom of a physical system, the resulting latent representation tends to align with these [iten2020dicvovering, nautrup2022operationally, fernandez2024learning].
II.2 VAE with a probabilistic decoder
@@
Figure 1: Schematic representation of the proposed pipeline. a) An experimental setup produces snapshots of a quantum system for different values of the experimental parameters $\theta_{1,2}$. b) A variational autoencoder (VAE) is trained on an ensemble of unlabeled snapshots collected from the previous experiment across a wide parameter range. Inspecting the learn latent neurons $\mathbf{z}$ across the parameter space will shed light about the system’s phase space (see Section III for more). c) Schematic representation of the cpVAE. The encoder neural network, ENC, takes as input a spin configuration. The latter outputs the means $\mu_{i}$ and the variances $\sigma_{i}$ parameterizing the Gaussian distribution from which the latent variables $z_{i}$ are sampled. Then, the decoder neural network DEC takes as input the latent variables and outputs the conditional probabilities of each spin be in a given state (up or down in the present scheme). During training, the decoder also receives as input the spin configuration to be reconstructed by the conditional probabilities, as indicated with the dotted arrow. Once trained, the latent neurons $z_{i}$ encode the main physical features of the physical system, which directly relate to its phase space. Moreover, one can fix the value of the latent variables and generate new configurations by autoregressively sampling the output conditional probabilities.
@@
As discussed in the introduction, prior applications of VAEs to quantum data have largely employed architectures with deterministic decoders [walker2020deep, wetzel2017unsupervised, jang2024unsupervised, naravane2023semi, frohnert2024explainable]. In these models, the decoder is trained to reconstruct individual spin configurations using a deterministic reconstruction loss between input and output spin configurations, as for instance the mean squared error (MSE). This approach implicitly assumes a deterministic relationship between the latent representation and the output configuration, neglecting the intrinsic probabilistic nature of quantum measurement outcomes. Moreover, the use of a deterministic reconstruction metric enforces the model to learn average configurations across samples rather than capturing the full probability distribution of outcomes. This leads to a latent space that reflects only the mean-field structure of the data, failing to capture any probabilistic feature, as for instance correlations between nearby spins. In what follows, we refer to this baseline model as the deterministic VAE (dVAE) to distinguish it from the stochastic decoder architecture introduced in this work.
@@
Motivated by the recent work on VAEs for stochastic processes [fernandez2024learning] and arguments on using generative classifiers for phase-classification [arnold2024mapping], we present here the key changes that will allow VAE to correctly create interpretable representations of quantum data. First, we construct a decoder inspired by the recent advances in Neural Quantum States (NQS) [lange2024from]. For example, NQS have recently demonstrated impressive performance in ground-state searches and non-equilibrium dynamics [chen2025convolutional]. By integrating an NQS-based decoder into the VAE framework, we ensure that the model can accurately capture and regenerate the statistical structure of the input quantum states.
@@
To account for the two previous arguments, we propose the use of an auto-regressive neural network (ARNN) for the decoder (see Fig. 1c and Appendix A for further details), and train it to model the conditional probabilities of each spin configuration, given the preceding spins. More precisely, the configuration of spin $i$ is given by the Bernoulli distribution. Its probability mass function $f$ over the two possible outcomes $x_{i}\in\{1,-1\}$ is
@@
To compute the KL regularization (second term in Eq. 1), we adopt the decomposition introduced in the Total Correlation Variational Autoencoder (TC-VAE) framework [chen2018isolating]. This approach enables a fine control over the structure of the latent space, by explicitly penalizing statistical dependencies between latent variables. This enforces the distinct degrees of freedom of some input physical data to be stored in independent latent neurons. The ELBO TC-decomposition consists in rewriting the KL regularization as
@@
In the latter, the first term on the right-hand side is the mutual information $I_{q}(z;x)$. The second term, referred to as the total correlation (TC), encourages the model to learn as so-called disentangled representations [chen2018isolating], where individual latent units capture distinct, minimally correlated physical features of the data. The third term is the dimension-wise KL which aligns the latent variables with the prior $\mathcal{N}(0,1)$. The hyperparameters, $\alpha$, $\beta$ and $\gamma$ are introduced to control the weight of each of these terms. The full loss with which the model is trained is thus
@@
$-\alpha I_{q}(z;x)-\beta\mathrm{KL}(q(z)||\prod_{j}q(z_{j}))$ (7)
@@ III Results
The first spin model is the transverse-field Ising model (TFIM) with next-nearest-neighbor (NNN) interaction. For weak NNN coupling, the model effectively reduces to the conventional TFIM, a setting where prior studies have applied deterministic VAEs [walker2020deep, wetzel2017unsupervised, jang2024unsupervised]. In this regime, we show that this success is largely due to the fact that the relevant order parameter is a mean-field observable. As the strength of the NNN interaction increases, the system deviates from such mean-field behavior, and we observe that the deterministic VAE fails to reconstruct physical features, in contrast to the probabilistic one, effectively hindering the representation of the system in their latent space.
@@
The code needed to reproduce the results presented in each figure is available in the accompanying code repository [github].
@@ III.1.1 NNN-TFIM
We start by considering the Transverse Field Ising Model with nearest neighbor and next-nearest-neighbor interaction (NNN-TFIM) in a one dimensional chain, whose Hamiltonian has the form[kassan2001one]
@@
with $J_{2}\in[0,1]$ and $X^{i},Z^{i}$ being the Pauli operators on spin $i$. The system under consideration has a size of $N=20$ with periodic boundary conditions.
The phase diagram of the NNN-TFIM can be characterized using two observables: the next-nearest-neighbor correlator $\langle Z^{i}Z^{i+2}\rangle$ and the magnetization $\langle Z^{i}\rangle$, as shown in Fig. 2a. For coupling strengths $J_{2}<0.5$, the nearest-neighbor interaction dominates, and the system effectively reduces to the conventional TFIM, a regime where deterministic VAEs had previous success [walker2020deep, wetzel2017unsupervised, jang2024unsupervised].
@@
We then train the cpVAE with configurations arising from different points in the phase space of this spin model. During training, we see that a single neuron remains active, while all other neurons collapse to the prior and are noised out (see Appendix B for further details). This is to be expected since, as shown in Fig. 2a, the whole phase space can be characterized only by the second nearest-neighbor correlator. To mitigate the effect of hyperparameter tuning and to make the comparison as fair as possible, we then train the dVAE without KL regularization and with a latent dimension of 1. By doing so, we substantially ease its training, and ensure that its latent representation is minimal. We note that not doing so, and considering a larger latent space, would allow the dVAE to store some stochastic features of the input sample in the latent space, for example, particular spins being up or down. While the KL term may help to find a more interpretable representation, our aim here is not to compare the structure of the latent representations, but rather to demonstrate that the dVAE cannot learn parts of the phase space due to its deterministic nature, in contrast to the cpVAE. A more detailed explanation of this benchmark strategy can be found in Appendix B.
@@
We then benchmark the generative capabilities of both models. To do so, we first input configurations from the whole phase space to the encoder, extract the corresponding latent space, and then sample new configurations from the decoder based on the extracted latent variables. In Fig. 2b, we show the $\langle Z^{i}Z^{i+2}\rangle$ correlator for the output configurations of the dVAE across the phase.
While the model accurately reproduces the expected correlations in both the ferromagnetic and paramagnetic regimes, it fails to capture the antiferromagnetic phase. Importantly, this regime cannot be characterized by the magnetization only, the usual order parameter extracted in previous works. Since the alternating bond structure inherent to the frustrated ground state falls beyond any mean-field parameter, the deterministic decoder is unable to correctly reproduce this statistics, and hence the antiferromagnetic phase will never appear in the latent representation of a dVAE.